Opened 8 years ago

Last modified 8 years ago

#404 assigned flight processing

GB11_01, flight day 165b/2011, Diddington

Reported by: knpa Owned by: knpa
Priority: whenever Milestone: 2011 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by emca)

Data location: ~arsf/arsf_data/2011/flight_data/uk/GB11_01-2011_165b_Diddington

Data arrived from ARSF via network transfer on 15/06/2011

GB11_01 Scientific objective:
Detection of Archaeological Residues using remote sensing Techniques (DART). DART will develop a deeper understanding of the contrast factors and detection dynamics, particularly
within traditionally ‘difficult’ areas (for example clay and pasture). This will allow the identification of
appropriate sensors and conditions for feature detection.

GB11_01 Priority: Whenever

PI: Doreen Boyd

PI: Anthony Beck

Notes: Pilot study - low priority.

Sensors:

Camera (13/07/2011)
Eagle (13/07/2011)
Hawk (13/07/2011)
LiDAR (15/11/2011)

Change History (18)

comment:1 Changed 8 years ago by knpa

  • Description modified (diff)

comment:2 Changed 8 years ago by emca

Starting Camera, Eagle and Hawk processing.

comment:3 Changed 8 years ago by emca

  • Owner set to emca
  • Status changed from new to assigned

comment:4 Changed 8 years ago by emca

No sync for eagle/hawk line 6

comment:5 Changed 8 years ago by emca

Eagle line 2 - no sync. Incorrect GPS start time in header file. Changed GPS start time in header file to match Hawk.

comment:6 Changed 8 years ago by emca

Hyperspectral data ready for delivery check

comment:7 Changed 8 years ago by emca

Camera ready for delivery check

comment:8 Changed 8 years ago by anch

Delivery checking hyperspectral

comment:9 Changed 8 years ago by anch

checking camera as well

comment:10 Changed 8 years ago by anch

Camera looks fine except perhaps "GB11_01" project code in readme is different to "GB11/01" in hyperspectral readme.

comment:11 Changed 8 years ago by anch

Project has been blocked because we're waiting for OS vectors

comment:12 Changed 8 years ago by emca

PI was happy to receive data without OS vector check. When the vectors arrive the data can be checked retrospectively.

Hyperspectral and Camera have been sent 13/07/2011. Disk was sent to David Stott @ Leeds as requested by PI.

comment:13 Changed 8 years ago by emca

Processed LiDAR data - ready for delivery check

comment:14 Changed 8 years ago by adbe

LiDAR delivery check:

  • There is not a logsheet
  • Maybe mention there is heavy cloud over some lines

Otherwise, ready to deliver

comment:15 Changed 8 years ago by emca

  • Description modified (diff)

Added logsheet to delivery. Data drop out was due the presence of a water body to east side of the survey area.

Data sent to David Stott @ Leeds as requested by PI 15/11/2011.

comment:16 Changed 8 years ago by mark1

Hyperspectral hawk bad pixel files and XML files have been re-created and re-delivered to PI (via email). This corrects sample number in bad pixel text files and Eagle IFOV value in XML and extra comments included.

comment:17 Changed 8 years ago by adbe

Hyperspectral sct's

lineeaglehawk
1-0.02-0.05
22.35-0.06
3-0.07-0.07
4-0.01-0.38
5-0.09-0.08
60.512.43
7-0.03-0.07
8-0.01-0.06

comment:18 Changed 8 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from emca to knpa

Beginning archiving.

Note: See TracTickets for help on using tickets.